OpenStudy (anonymous):

Write out the form of the partial fraction decomposition of the function (See Example). Do not determine the numerical values of the coefficients. (If the partial fraction decomposition does not exist, enter DNE.) (a) (x)/(x^2 + x − 6)

hartnn (hartnn):

can you factor x^2+x-6 ?

OpenStudy (anonymous):

(x+3) (x-2)

hartnn (hartnn):

good, so for \(\huge \dfrac{x}{(x-a)(x-b)}\) it'll be \(\huge \dfrac{A}{x-a}+\dfrac{B}{x-b}\) so what about \(\huge \dfrac{x}{(x+3)(x-2)}\)

hartnn (hartnn):

still there ?